# A mechanism-informed reinforcement learning framework for shape optimization of airfoils

### v1

@@ 3 Automatic DWR-based adaptation for the target functional
where $\mathbf{u}_{h}^{H}$ denotes the coarse mesh solution $\mathbf{u}_{H}$ projected onto the fine mesh space $\mathcal{V}_{h}$ through a prolongation operator $I_{h}^{H}$,
$\mathbf{u}_{h}^{H}=I_{h}^{H}\mathbf{u}_{H}.$ (8)
We identify the residual of primal problem in the space $\mathcal{V}_{h}$ by $\mathcal{R}_{h}(\cdot)$, which is expected to satisfy the equations with solution $\mathbf{u}_{h}$, i.e.,
$\mathcal{R}_{h}(\mathbf{u}_{h})=0.$ (9)
Upon linearizing, we derive the following expression,
@@
$\mathbf{u}_{h}-\mathbf{u}_{h}^{H}\approx-\left(\left.\frac{\partial\mathcal{R}_{h}}{\partial\mathbf{u}_{h}}\right|_{\mathbf{u}_{h}^{H}}\right)^{-1}\mathcal{R}_{h}(\mathbf{u}_{h}^{H}).$ (11)
@@
The calculation of $\mathbf{z}_{h}$ on the embedded mesh was normally expensive. The literature presents various methods to streamline this process. For instance, the works of [47, 49] introduce an error correction method that interpolates dual solutions onto a fine mesh, while [50] proposes a smoothing technique to mitigate oscillations stemming from interpolation. In recent years, the BiCG method in [10, 51] has been adopted to calculate the dual and primal problems at once. With the help of machine learning techniques, we developed the hybrid CNNs-Dual framework[33] to accelerate the calculation of dual solutions which is very effective. However, the CNNs-Dual solver can not preserve the accuracy for deformed geometry. While techniques like GCN(Graph Convolutional Networks)-enhanced learning should be developed to mend this issue, a conventional algorithm for the DWR-based adaptation is adopted here to calculate the target functional.
Even though the dual solutions can be solved effectively, issues persist on how to balance the value of primal residuals and dual solutions as their magnitudes are not compatible. Indicators for each element are generated based on their multiplication, i.e.,
$\eta_{K_{i}}=\sum_{K_{ij}}|(\mathbf{z}_{h})^{T}\mathcal{R}_{h}(\mathbf{u}_{h}^{H})|,$ (14)
@@
where $K_{ij}$ is the subelements $K_{i}$. For the implementation of mesh adaptation, a parameter needs to be set for guiding such a process. However, setting this parameter is tricky because a lower value may lead to an approximately uniform refinement while a higher value may not solve the target functional with expected accuracy. What makes it worse is that such a parameter differs with the update of configurations and different geometries. Thus, it is crucial to find an automatic parameter-setting strategy for guaranteeing the mesh adaptation process during the optimization. In [33], we developed such a technique based on statistical analysis. At last, the procedure of the algorithm for calculating the target functional is concisely recapitulated below:
Data: Initial mesh $\mathcal{K}_{H}$
Result: Refined mesh $\mathcal{K}_{h}$ and target functional $J_{h}(\mathbf{u}_{h})$
1 while Iter < refine_step do
2       Set the current mesh as $\mathcal{K}_{H}$;
3       Solve $\mathcal{R}_{H}(\mathbf{u}_{H})=0$ with a residual tolerance of $1.0\times 10^{-3}$;
4       Interpolate solution $\mathbf{u}_{H}$ from mesh $\mathcal{K}_{H}$ to $\mathcal{K}_{h}$ to obtain $\mathbf{u}_{h}^{H}$;
5       Compute the residual $\mathcal{R}_{h}\left(\mathbf{u}_{h}^{H}\right)$;
6       Solve $\mathcal{R}_{h}(\mathbf{u}_{h})=0$ with a residual tolerance of $1.0\times 10^{-3}$;
7       Generate the dual solutions from the primal residuals;
8       Obtain $\mathbf{z}_{h}$ by solving the dual equation with GMG;
9       Determine the error indicator for each element;
10       Automatic strategy for selecting $TOL$;
11       while any element $K_{H}$ has $\mathcal{E}_{K_{H}}>TOL$ do
12             Refine element $K_{H}$ adaptively as described in [31];
13       end while
14
15 end while
16Obtain the target functional $J_{h}(\mathbf{u}_{h})$;
@@ 12             Refine element $K_{H}$ adaptively as described in [31];
This efficient algorithm ensures the accuracy of the target functional is maintained across configurations and geometries not typically encountered in engineering problems. However, during the exploration process inherent in the reinforcement learning framework, scenarios may arise that deviate significantly from those within the training datasets. Under these circumstances, a learning-based solver generally struggles to maintain accuracy, potentially feeding incorrect experiences back to the agent. In contrast, the paradigm of optimizing via reinforcement learning based on PDEs offers a more precise framework to assist the agent in making judgments.
It’s important to highlight that the DWR-based adaptation method cannot be universally applied to all types of target functionals. Specifically, this analysis does not extend to non-smooth versions. For generalized target functionals, such as the lift-drag ratio, a multi-mesh method, as developed in [52, 53], should be employed. In this study, our focus is primarily on the drag of airfoils, which exhibit superior robustness according to our previous study.
4 A reinforcement learning framework
@@ 4.1 Markov decision process
The iteration process of reinforcement learning can be treated as an agent interacting with the environment and accumulating experience from the actions as illustrated in Figure 3. The shape of the wing, composed of hundreds of points, can be conceptualized as a vector in a high-dimensional space, essentially representing a point in this space. Thus, all potential wing shapes can be understood as forming a submanifold within this high-dimensional space. The objective function, in this context, is a functional defined on this submanifold, and the goal of the reinforcement learning process is to locate the extremum of this functional on the submanifold. To invoke such an agent for developing the expected airfoil, the optimization process should be modeled as the Markov decision process(MDP) at first.
Figure 3: Reinforcement Learning Framework.
@@
At each time step $t$, the standard MDP process is defined by a 4-tuple $(S_{t},A_{t},P_{t},R_{t})$ [54] which contains the interaction between agent and environment, where $t=0,1,2,\cdots.$ Here $S_{t}$ represents the state, which consists of the mesh points that depict the geometry of the airfoil at time step $t$. $A_{t}$ stands for the action performed on $S_{t}$. The action $A_{t}$ consists of three parameters $[x_{target},y_{Uchange},y_{Lchange}]$. For a given $x$ coordinate $x_{target}$, there exists a pair of points on both curves whose values on $y$ direction are $y_{upper}$ and $y_{lower}$. Then the $y$ coordinates will be updated with parameters $y_{Uchange}$ and $y_{Lchange}$ on both curves respectively. However, the values will not only be acted with respect to $x_{target}$. This behavior introduces a sharp deformation, which will bring singularity to the geometry and may even cause a mesh tangling phenomenon. To perform a more smooth and natural action on the airfoil, we utilize a smooth technique as shown in Figure 4. For a specific $x_{target}$, the update value $y_{Uchange}$ will not be acted on the point directly. The delta function is defined as
$SA(x)=e^{-\frac{(x-x_{target})^{2}}{2\delta^{2}}}$ (15)
will be set as a coefficient that can smoothen the deformation process. $\delta$ is set to $0.2$ to $0.8$ in this model. Then the $y$ coordinate of the upper curve will be updated with
$y^{\prime}=y+y_{Uchange}SA(x).$ (16)
@@
Figure 4: Update curve with a smooth action.
As explained above, a specific airfoil shape is a single point in a high-dimensional space. If the update value is just an entry in the vector formulated by shape, the difference between the update vector and the original vector is weeny. As a result, the agent may not notice such discrepancy and is not capable of performing tactful action. The smoothing process not only improves the efficiency during the optimization but also assists the action learning process by amplifying the difference between tensors.
@@
After the deformation, the algorithm of curve fitting will generate the Bézier curves, which will be explained later. The update points represent the new state of the airfoil, $S_{t+1}$. At each time step, there are various actions that the agent can take. $p(S_{t+1}|S_{t})$ is denoted as the transition probability for the state $S_{t}$ change to $S_{t+1}$ with action $A_{t}$. The agent will evaluate the potential reward from different actions. The action that the agent takes at state $S_{t}$ is $A_{t}$, which is generated from the policy $\pi_{t}:S_{t}\rightarrow A_{t}$. The agent will evaluate the reward $R_{t+1}(S_{t},A_{t},S_{t+1})$ from the environment and try to maximize the cumulative reward from a potential trajectory $\tau$, which is given by
$\tau=(S_{0},A_{0},S_{1},A_{1},\cdots,S_{t},A_{t},\cdots).$ (17)
In this trajectory, the cumulative reward at time step $t$ is defined as
$G_{t}(\tau)=\sum_{i=0}^{T}\gamma^{i}R_{t+i+1}.$ (18)
Then the agent shall be trained to maximize the reward so that the deformation can fulfill the optimization target. Various deep learning algorithms have been constructed in the past decades like the Deep Q Networks(DQN), Double Deep Q Networks (DDQN)[55], Proximal Policy Optimization[56] (PPO) and Twin Delayed DDPG(Deep deterministic policy gradient)[57]. Among the algorithms, the Twin Delayed DDPG algorithm is constructed in this work for its efficiency in continuous space and is suitable for the current framework. In the following subsection, we start with the original DDPG algorithm to understand the shape optimization process with reinforcement learning. Then the Twin Delayed DDPG algorithm is a melioration based on this framework.
4.2 Deep deterministic policy gradient for training
@@
We start with the basic DDPG algorithm to understand the optimization process of the airfoil shape optimal design. The DDPG framework consists of two primary components: the actor and the critic, each with its distinct role in learning the optimal policy. The actor, parameterized by $\theta^{a}$, is responsible for determining the best possible action $A_{t}$ given the current state $S_{t}$ at time step $t$ with the current policy, i.e., $A_{t}=\pi_{\theta^{a}}(S_{t})$. This direct policy optimization allows the DDPG algorithm to efficiently explore the continuous action space and identify optimal actions.
@@
Since the actor may generate potential actions that form a state-action sequence, i.e., $\tau_{\theta^{a}}\sim\pi_{\theta^{a}}$. The cumulative reward at time step $t$ is not explicitly calculated because the reward should be obtained from the CFD solver, which is generally time-consuming. The critic, on the other hand, parameterized by $\theta^{c}$ evaluates the quality of the actions proposed by the actor within the current policy. Hence, $\tilde{R}_{t+1}(S_{t},A_{t}|\theta^{c})$ is the reward calculated by the critic with a neural network. For a specific long term state evolution $\tau$,
$\tilde{G}_{t}(\tau_{\theta^{a}}|\theta^{c})=\tilde{R}_{t+1}(S_{t},A_{t}|\theta^{c})+\sum_{i=1}^{T}\gamma^{i}\tilde{R}_{t+i+1}\Big{(}(S_{t+i}|\theta^{a}),(A_{t+i}|\theta^{a})|\theta^{c}\Big{)}.$ (19)
@@
$Q(S_{t}|\theta^{a},\theta^{c})=\mathbb{E}_{\tau_{\theta^{a}}\sim\pi_{\theta^{a}}}\Big{(}\tilde{G}_{t}(\tau_{\theta^{a}}|\theta^{c})\Big{)}$ (20)
is adopted to assess the performance of the actor with current parameters $\theta^{a}$. The critic guides the actor towards actions that maximize the long-term reward. The feedback of critic is essential for updating the parameters of actor and refining the policy.
During training, a replay buffer is utilized to store and randomly sample previous transition tuples $(S_{t},A_{t},S_{t+1},R_{t+1})$ facilitating the learning of uncorrelated and diversified experiences. This experience replay mechanism significantly stabilizes and enhances the training process by preventing the correlation of consecutive experiences and reducing the variance of the updates.
In the training process, the critic is updated by minimizing the mean squared error between a batch of current Q-values and the batch of target Q-values, which are calculated based on the next state and the next action proposed by the actor. The actor is then updated by performing gradient ascent on the expected Q-values output by the critic, thus steering the actor towards actions that yield higher rewards.
Through the interaction between the actor and the critic, the DDPG framework effectively navigates the continuous action space with the experience in the replay buffer. As a result, the agent is able to learn sophisticated policies and perform expected actions on the airfoil, thereby optimizing its performance in complex environments. A brief algorithm of this DDPG training process is organized as follows:
Data: Wing shape data, DDPG agent with actor and critic networks, ReplayBuffer
Result: Trained DDPG agent
1 for each episode in max_episodes do
2       Reset the environment and get the initial state;
3       while the episode is not done do
4             The agent selects and executes an action;
5             Observe the new state and the reward;
6             Save the tuple (state, action, reward, next_state) to the ReplayBuffer;
7             if ReplayBuffer has enough experiences then
8                   Sample a batch of experiences from ReplayBuffer;
9                   Use the batch of experiences to train the actor and critic networks;
10
11             end if
12            Update the state;
13
14       end while
15
16 end for
Algorithm 2 Deep Deterministic Policy Gradient (DDPG)
4.3 Mesh deformation for unstructured mesh
The mesh deformation during the shape optimization is a complicated process. It is important to ensure the quality of the mesh is capable of the calculation of target functional. The process of mesh deformation, modification, and calculation process can be organized as shown in Figure 5.
@@ 4.3 Mesh deformation for unstructured mesh
Figure 6: Airfoil consists of two parts.
With the Bézier least square fitting, a series of control points can be calculated with the mesh points along the airfoil where the number of control points can be set manually. Then the Bézier curves can be generated by the Bernstein polynomials, i.e.,
@@
Here $B(t)$ represents the Bézier curve itself, which is a function of the parameter $t$. $t$ ranges from 00 to $1$ which essentially interpolates between the control points $CP_{i}$. Then we set the airfoil with the defined curve in a circle with a large enough radius. With such a technique, we defined the computational domain on an unstructured mesh, and the mesh around the airfoil is shown in Figure 1. The airfoil is depicted with a determinate number of mesh points. The order of these mesh points is arranged in an anticlockwise direction from the tail endpoint and forms a closed cycle.
When the agent of reinforcement learning tries to modify the shape of airfoils, it will change all the mesh points. Equivalently, the control points will be updated and result in varied Bézier curves. From the updated expression of Bézier curves, we can calculate the mesh points along the airfoils in an anticlockwise direction. Afterward, we build a map structure so that the mesh file can find the corresponding mesh points on the airfoils and update their coordinates. As shown in the left part of Figure 7, the geometry received the action from the agent and the shape gets updated.
@@
Figure 7: Left: mesh around the deformed shape before the modifying; Right: mesh around the deformed shape after the modifying;
@@
However, it is shown in the left part of Figure 7 that the quality of mesh gets influenced since the movement of boundary mesh points may lead to a sharp acute angle for the triangles near the upper curve and a large obtuse angle for the triangles near the lower half. While the optimization lasts for a long period, the phenomenon of mesh tangling will get more serious with iteration steps. Hence we shall modify the mesh so that it is capable of multi-step mesh deformation. In numerical simulation, it is never suggested to bring sharp transformation to the numerical scheme like geometry and solution update. Hence the movement of points is restricted to a specific scale. Then the deformed mesh should be modified for further calculation. Since the most closely affected elements are those on the boundary part, such area should be processed specially. For those boundary elements that have two points on the airfoil, the third point is relocated so that this element will be a regular triangle. Then we adopted the Laplacian smoothing technique to amend the quality of the mesh. The fundamental operation in Laplacian smoothing involves updating the position of a node to be the average of the positions of its neighboring nodes. Mathematically, this can be expressed as
$\mathbf{x}_{i}^{{}^{\prime}}=\frac{1}{m}\sum_{j=1}^{m}\mathbf{x}_{j},$ (22)
where $\mathbf{x}_{i}$ is the coordinate vector of the $i$-th point and $\mathbf{x}_{i}^{{}^{\prime}}$ is the update value. $\mathbf{x}_{j}$ with index from $1$ to $m$ are neighboring nodes of corresponding $i$-th point. With such a modification process, the quality of the mesh is enhanced. As shown in the right part of Figure 7, the elements around the airfoil are recovered so that sharp acute angle and large obtuse angle can be prevented.
@@
To implement such mesh recovery within our algorithm, only adopting smoothing techniques is not sufficient as it cannot preserve the desired properties of boundary-sensitive points. Aiming to maintain efficiency throughout the iterative process, remeshing is not considered since it is time-consuming, especially when generalized to three-dimensional cases. Even though the post-smoothing mesh possesses improved structural integrity of the triangular elements, it should not be far away from the pre-processed state. To this end, we utilize the map data structure inherent in C++ for its capability to manage modifications without drastically altering point positions or changing the number of vertices within the mesh.
@@
In our approach, vertices that are not on the boundary but belong to boundary triangles are assigned with level 1. Subsequently, vertices without a level, belonging to triangles that include level 1 vertices, are marked as level 2. We then recalculate the triangles containing level 1 and level 2 vertices so that the new triangles can approximate an equilateral triangle as possible. This recalculated position is mapped against the initial position, facilitating updates to the mesh data file through these mappings. Thus, without a remeshing process, we achieve a recovery of mesh properties. It helps the mesh remain similar to its original state while significantly enhancing the quality of the triangular elements. This method allows for an effective improvement in mesh quality, crucial for the accuracy of subsequent simulations and analyses.
@@
It is worth noting that when the agent tries to modify the geometry, the curvature of the airfoil changes rapidly. The deformation conducted with a scale comparable to the mesh size can be handled with the method mentioned above. However, when the scale is smaller than the mesh size, the given mesh points can not detect such curvature variation within the segment. Then, it will add noise to both the primal and dual equations, and return an imprecise reward back to the agent. As a result, it will bring more complexities and increase training difficulties. For this reason, we shall use triangles with a smaller scale.

### v2

@@ 2.2 An automatic DWR-based adaptation for the target functional calculation
where $\mathbf{u}_{h}^{H}$ denotes the coarse mesh solution $\mathbf{u}_{H}$ projected onto the fine mesh space $\mathcal{V}_{h}$ through a prolongation operator $I_{h}^{H}$, $\mathbf{u}_{h}^{H}=I_{h}^{H}\mathbf{u}_{H}.$
We denote the residual of governing equations in the space $\mathcal{V}_{h}$ by $\mathcal{R}_{h}(\cdot)$, which satisfies the equations $\mathcal{R}_{h}(\mathbf{u}_{h})=0.$ Upon linearizing, we derive
@@
$\mathbf{u}_{h}-\mathbf{u}_{h}^{H}\approx-\left.({\partial\mathcal{R}_{h}}/{\partial\mathbf{u}_{h}})^{-1}\right|_{\mathbf{u}_{h}^{H}}\mathcal{R}_{h}(\mathbf{u}_{h}^{H}).$ (10)
@@
The calculation of $\mathbf{z}_{h}$ on the embedded mesh was expensive. The literature presents various methods to streamline this process. For instance, the works of [41, 43] introduce an error correction method that interpolates dual solutions onto a fine mesh, while the BiCG method in [8, 44] has been adopted to calculate the dual and primal problems at once. With the help of machine learning techniques, we developed the hybrid CNNs-Dual framework[32] to accelerate the calculation of dual solutions which is very effective. However, the CNNs-Dual solver can not preserve the accuracy for deformed geometry. Even though techniques like GCN(Graph Convolutional Networks)-enhanced learning can mend this issue, we still adopt the DWR-based adaptation developed in [31] in this work for its robustness.
Refer to captionRefer to caption
@@
Although the dual solutions can be solved effectively, issues persist in balancing the value of primal residuals and dual solutions as their magnitudes are not compatible. Indicators for each element are generated based on their multiplication $\eta_{K_{i}}=\sum_{K_{ij}}|(\mathbf{z}_{h})^{T}\mathcal{R}_{h}(\mathbf{u}_{h}^{H})|,$ where $K_{ij}$ is the subelements $K_{i}$. For the implementation of mesh adaptation, a parameter needs to be set for guiding such a process. However, setting this parameter is tricky because a lower value may lead to an approximately uniform refinement while a higher value may not solve the target functional with expected accuracy. What makes it worse is that such a parameter differs with the update of configurations and different geometries. Thus, it is crucial to find an automatic parameter-setting strategy for guaranteeing the mesh adaptation process during the optimization. In [32], we developed such a technique based on statistical analysis. At last, the procedure of the algorithm for calculating the target functional is concisely organised as algorithm 1:
0:  Initial mesh $\mathcal{K}_{H}$
0:  Refined mesh $\mathcal{K}_{h}$ and target functional $J_{h}(\mathbf{u}_{h})$
1:  while Iter $<$ refine_step do
2:     Set the current mesh as $\mathcal{K}_{H}$
3:     Solve $\mathcal{R}_{H}(\mathbf{u}_{H})=0$ with a residual tolerance of $1.0\times 10^{-3}$
4:     Interpolate solution $\mathbf{u}_{H}$ from mesh $\mathcal{K}_{H}$ to $\mathcal{K}_{h}$ to obtain $\mathbf{u}_{h}^{H}$
5:     Compute the residual $\mathcal{R}_{h}\left(\mathbf{u}_{h}^{H}\right)$
6:     Solve $\mathcal{R}_{h}(\mathbf{u}_{h})=0$ with a residual tolerance of $1.0\times 10^{-3}$
7:     Generate the dual solutions from the primal residuals
8:     Obtain $\mathbf{z}_{h}$ by solving the dual equation with GMG
9:     Determine the error indicator for each element
10:     Automatic strategy for selecting $TOL$
11:     while any element $K_{H}$ has $\mathcal{E}_{K_{H}}>TOL$ do
12:        Refine element $K_{H}$ adaptively as described in [30]
13:     end while
14:  end while
15:  Obtain the target functional $J_{h}(\mathbf{u}_{h})$
@@
This efficient algorithm ensures the accuracy of the target functional can be maintained across configurations and geometries not typically encountered in engineering problems. During the exploration process inherent in the reinforcement learning framework, scenarios may arise that deviate significantly from those within the training datasets. Under these circumstances, a learning-based solver generally struggles to maintain accuracy, potentially feeding incorrect experiences back to the agent. In contrast, the paradigm of PDEs-constrained optimizing via reinforcement offers a more precise framework to assist the agent in making judgments.
3 A reinforcement learning framework for shape optimization of airfoils
The iteration process of reinforcement learning can be treated as an agent interacting with the environment and accumulating experience. The shape of the airfoil, composed of hundreds of points, can be conceptualized as a vector in a high-dimensional space, essentially representing a point in this space. Thus, all potential airfoil shapes can form a set within this high-dimensional space. The objective function, in this context, is a functional defined on this set. The goal of the reinforcement learning process is to locate the extremum on the set. To discuss the learning framework for the shape optimization of airfoil, the Markov decision process(MDP) with specific components should be defined correspondingly. However, the mesh deformation conducted within the unstructured mesh is challenging, where mesh tangling, curvature oscillation and irregular shapes are potential issues that will greatly influence the learning process. In this section, we will introduce a basic process of the reinforcement learning and address these potential issues.
@@ 3.1 Markov decision process
The Markov decision process can be defined as follows. At each time step $t$, a 4-tuple $(S_{t},A_{t},P_{t},R_{t})$ is adopted to record the process [45] which contains the interaction between agent and environment, where $t=0,1,2,\cdots.$ Here $S_{t}$ represents the state, consisting of the mesh points that depict the geometry of the airfoil at time step $t$. $A_{t}$ stands for the action performed on $S_{t}$. The action $A_{t}$ consists of three parameters $[x_{target},y_{Uchange},y_{Lchange}]$. Given $x$ coordinate $x_{target}$, there exists a pair of points on both curves whose values on $y$ direction are $y_{upper}$ and $y_{lower}$. Then the $y$ coordinates will be updated with parameters $y_{Uchange}$ and $y_{Lchange}$ on both curves respectively. However, the values will not only be acted with respect to $x_{target}$. This behavior introduces a sharp deformation, which will bring irregularity to the geometry and may even cause a mesh tangling phenomenon. To perform a more smooth and natural action on the airfoil, we utilize a smooth technique as shown in Figure 2. There are totally 132 sample points to depict the shape of the airfoil. For a specific $x_{target}$, the update value $y_{Uchange}$ will not be acted on the point directly. The delta function, defined as $SA(x)=e^{-(x-x_{target})^{2}/2\delta^{2}},$ will be set as a coefficient that can act on the neighbor sample points and smoothen the deformation process. $\delta$ is set as $0.2$ to $0.8$ in this model. Then the $y$ coordinate of the upper curve will be updated with $y^{\prime}=y+y_{Uchange}SA(x).$
@@
Figure 2: Update action with a smooth function acted on the geometry with $132$ sample points.
@@
After the deformation, our algorithm will generate the Bézier curves. The update points represent the new state of the airfoil, $S_{t+1}$. At each time step, there are various actions that the agent can take. $p(S_{t+1}|S_{t})$ is denoted as the transition probability for the state $S_{t}$ change to $S_{t+1}$ with action $A_{t}$. The agent will evaluate the potential reward from different actions. The action that the agent takes at state $S_{t}$ is $A_{t}$, which is generated from the policy $\pi_{t}:S_{t}\rightarrow A_{t}$. The agent will evaluate the reward $R_{t+1}(S_{t},A_{t},S_{t+1})$ from the environment and try to maximize the cumulative reward from a potential trajectory $\tau:=(S_{0},A_{0},S_{1},A_{1},\cdots,S_{t},A_{t},\cdots).$ In this trajectory, the cumulative reward at time step $t$ is defined as $G_{t}(\tau)=\sum_{i=0}^{T}\gamma^{i}R_{t+i+1}.$ Then the agent shall be trained to maximize the cumulative reward so that the deformation can fulfill the optimization target. The Twin Delayed DDPG algorithm[46] is constructed in this work for its efficiency in continuous space and is suitable for the current framework. In the following part, we start with two important components during the learning process to understand such a learning task, actor and critic.
3.2 An actor-critic scheme
@@
The actor, parameterized by $\theta^{a}$, is responsible for determining the best possible action $A_{t}$, given the current state $S_{t}$ at time step $t$ with the current policy, i.e., $A_{t}=\pi_{\theta^{a}}(S_{t})$. This direct policy optimization allows the agent to efficiently explore the continuous action space and identify optimal actions.
@@
The actor will generate potential actions that form a state-action sequence, i.e., $\tau_{\theta^{a}}\sim\pi_{\theta^{a}}$. The cumulative reward at time step $t$ cannot explicitly calculated as the reward should be obtained from the CFD solver, which is generally time-consuming while the infinite sequence cannot be obtained. The critic, parameterized by $\theta^{c}$, evaluates cumulative reward based on corresponding states. Hence, $\tilde{R}_{t+1}(S_{t},A_{t}|\theta^{c})$ is the cumulative reward calculated by the critic with a neural network. For a specific long-term state evolution $\tau$,
$\tilde{G}_{t}(\tau_{\theta^{a}}|\theta^{c})=\tilde{R}_{t+1}(S_{t},A_{t}|\theta^{c})+\sum_{i=1}^{T}\gamma^{i}\tilde{R}_{t+i+1}\Big{(}(S_{t+i}|\theta^{a}),(A_{t+i}|\theta^{a})|\theta^{c}\Big{)}$ (13)
@@
$Q(S_{t}|\theta^{a},\theta^{c})=\mathbb{E}_{\tau_{\theta^{a}}\sim\pi_{\theta^{a}}}\Big{(}\tilde{G}_{t}(\tau_{\theta^{a}}|\theta^{c})\Big{)},$ (14)
is adopted to assess the performance of the actor based on $\theta^{a}$. The critic guides the actor towards actions that maximize the long-term reward. The feedback of critic is essential for updating the parameters of actor and refining the policy.
During training, a replay buffer is utilized to store and randomly sample previous transition tuples $(S_{t},A_{t},S_{t+1},R_{t+1})$ , facilitating the learning of uncorrelated and diversified experiences. In the training process, the critic is updated by minimizing the mean squared error between a batch of current Q-values and target Q-values. The actor is then updated by performing gradient ascent on the expected Q-values output by the critic, thus steering the actor towards actions that yield higher rewards.
Through the interaction between the actor and the critic, the agent effectively navigates the continuous action space with the experience in the replay buffer. As a result, it can learn sophisticated policies and perform expected actions on the airfoil, thereby optimizing its performance in complex environments. However, performing actions on this problem, based on such a high-dimensional parameters within the unstructured mesh, is not straightforward. Further improvement should be conducted.
3.3 Action of mesh deformation on unstructured mesh
Mesh deformation during shape optimization is a complicated process. It is important to ensure the quality of the mesh is capable of calculating target functional. The process of mesh deformation, modification, and calculation can be organized as Figure 3.
@@ 3.3 Action of mesh deformation on unstructured mesh
With the Bézier least square fitting, control points can be calculated with the sample points along the airfoil. Then the Bézier curves can be generated by the Bernstein polynomials, i.e.,
@@
Here $B(t)$ represents the Bézier curve, which is a function of the parameter $t$. $t$ ranges from 00 to $1$ which essentially interpolates between the control points $CP_{i}$. Then the airfoil with the defined curve is set in a circle with a large enough radius. With such a technique, we defined the computational domain on an unstructured mesh. The airfoil is depicted with a determinate number of mesh points. However, difficulties will brought by the unstructured mesh and high dimensionality with the iteration get processed. We will elaborate these issues as follows.
•  Mesh tangling. When the agent tries to modify the shape of airfoils, it will change all the sample points. Consequently, the control points will be updated and result in varied Bézier curves. From the updated expression of Bézier curves, we can calculate the mesh points along the airfoils in an anticlockwise direction. Afterward, we build a map structure so that the mesh file can find the corresponding mesh points on the airfoils and update their coordinates. As shown in the left part of Figure 4, the geometry received the action from the agent and the shape gets updated.
@@
Figure 4: Meshes around the airfoils. Left: before modifying; Right: after modifying;
@@
However, it is shown in the left part of Figure 4 that the quality of mesh gets influenced since the movement of boundary mesh points may lead to a sharp acute angle for the triangles near the upper curve and a large obtuse angle for the triangles near the lower half. While the optimization lasts for a long period, the elements around the airfoil will become more irregular with iteration steps, leading to the mesh tangling issues. Hence we shall modify the mesh so that it is capable of dealing multi-step mesh deformation. The movement of points is restricted to a specific scale. Then the deformed mesh should be modified for further calculation. Since the most closely affected elements are those on the boundary part, such an area should be processed specially. For those boundary elements that have two points on the airfoil, the third point is relocated so that this element will be a regular triangle. Then we adopted the Laplacian smoothing technique to amend the quality of the mesh. The fundamental operation in Laplacian smoothing involves updating the position of a node to be the average of the positions of its neighboring nodes. Mathematically, this can be expressed as $\mathbf{x}_{i}^{{}^{\prime}}=\frac{1}{m}\sum_{j=1}^{m}\mathbf{x}_{j}$, where $\mathbf{x}_{i}$ is the coordinate vector of the $i$-th point and $\mathbf{x}_{i}^{{}^{\prime}}$ is the update value. $\mathbf{x}_{j}$ with index from $1$ to $m$ are neighboring nodes of the original point. With such a modification process, the quality of the mesh is enhanced. As shown in the right part of Figure 4, the elements around the airfoil are recovered so that sharp acute angle and large obtuse angle can be prevented.
@@
Aiming to maintain efficiency throughout the iterative process, re-meshing is not considered as it is time-consuming, especially when generalized to three-dimensional cases. Even though the post-smoothing mesh possesses improved structural integrity of the triangular elements, it should not be far away from the pre-processed state. To this end, we leverage the C++ Standard Template Library’s map data structure for its capability to manage modifications without drastically altering point positions or changing the number of vertices within the mesh.
@@
In our approach, vertices that are not on the boundary but belong to boundary triangles are assigned with level 1. Subsequently, vertices without a level, belonging to triangles that include level 1 vertices, are marked as level 2. Then we recalculate the triangles containing level 1 and level 2 vertices so that the new triangles can approximate an equilateral triangle as possible. This recalculated position is mapped from the initial position, facilitating updates to the mesh data file through these mappings. Thus, without a re-meshing process, it allows for an effective improvement in mesh quality, crucial for the accuracy of subsequent simulations and analyses.
@@
• Curvature oscillation. It is worth noting that when the agent tries to modify the geometry, the curvature of the airfoil changes rapidly. The deformation conducted with a scale comparable to the mesh size can be handled with the method mentioned above. However, when the scale is smaller than the mesh size, the given mesh points can not detect such curvature variation within the segment. Then, it will add noise to both the primal and dual equations, and return an imprecise reward back to the agent. As a result, it will bring more complexities and increase training difficulties. Out of this reason, we shall use triangles with a smaller scale.